the area of the pool is x squared plus 16x plus 48.what are the dimensions?
Answer by checkley71(8403)

AFTER SEEING THE ANSWERS iI QUESTION THE EQUATION?????????
x^2+16x+48=0
(x+12)(x+4)=0
x+12=0
x=-12
x=4=0
x=-4
THIS DOESN'T COMPUTE AS NEGATIVE DIMENTIONS AREN'T POSSIBLE.